# A Survey of Mamba

### v3

@@ 1. Introduction
Recently, a promising architecture, structured state space sequence models (SSMs) (Gu et al., 2021a), have emerged to efficiently capture complex dependencies in sequential data, becoming a formidable competitor to Transformer. These models, inspired by classical state space models (Kalman, 1960), can be considered a fusion of recurrent neural networks and convolutional neural networks. They can be computed efficiently using either recurrence or convolution operations, achieving linear or near-linear scaling with sequence length, thus significantly reducing the computational costs. More specifically, as one of the most successful SSM variants, Mamba achieves comparable modeling capabilities to Transformers while maintaining linear scalability with sequence length (Gu and Dao, 2023), propelling it into the realm of focal topics. Mamba first introduces a simple yet effective selection mechanism that enables the model to filter out irrelevant information while retaining necessary and relevant data indefinitely by parameterizing the SSM parameters based on the input. Then, Mamba proposes a hardware-aware algorithm to compute the model recurrently with a scan instead of convolution, achieving up to 3$\times$×faster computation on A100 GPUs. As shown in Figure 1, the powerful modeling capabilities for complex and lengthy sequential data, along with near-linear scalability, position Mamba as an emerging foundation model, poised to revolutionize various domains of research and applications, such as computer vision (Xu et al., 2024b; Zhu et al., 2024), natural language processing (Lieber et al., 2024; Zhao et al., 2024c), healthcare (Ruan and Xiang, 2024; Xing et al., 2024; Wang et al., 2024g), etc. For example, Zhu et al. (2024) propose Vim, which is 2.8$\times$×faster than DeiT (Touvron et al., 2021) and saves 86.8% GPU memory when extracting features for high-resolution images. Dao and Gu (2024) show the connections between SSMs and variants of attention and propose a new architecture that refines selective SSM, achieving 2-8$\times$ faster on language modeling.
@@
Concurrent with our survey, several related surveys have been released, purely focusing on state space models (Patro and Agneeswaran, 2024; Wang et al., 2024f) and Vision Mamba (Zhang et al., 2024h; Liu et al., 2024e; Xu et al., 2024b). Diverging from these surveys, this paper is centered on the associated research concerning Mamba. It systematically analyzes existing literature from a novel standpoint to explore the evolution of Mamba architecture and the data adaptation methods utilized in Mamba-based models.
@@ 3.2. Mamba-2: State Space Duality
(24) $\mathbf{y}=\text{SSD}(\mathbf{A},\mathbf{B},\mathbf{C})(\mathbf{x})=\mathbf{M}\mathbf{x},$
@@ 4. Advancements in Mamba Models
Block Integration Transformer (Lieber et al., 2024; Xu et al., 2024a; Pilault et al., 2024; Hatamizadeh and Kautz, 2024; Pitorro et al., 2024; Gao et al., 2024b)
@@
Graph Neural Network (GNN) (Liu et al., 2024a; Li et al., 2024f; Behrouz and Hashemi, 2024; Wang et al., 2024e; Yang et al., 2024d)
@@
Substitution U-Net (Sepehri et al., 2024; Shi et al., 2024b; Wang et al., 2024g, a; Liu et al., 2024d; Ruan and Xiang, 2024; Liao et al., 2024; Ma and Wang, 2024; Sanjid et al., 2024; Deng and Gu, 2024; Ji et al., 2024; Hosseini et al., 2024)
Diffusion Models (Oshima et al., 2024; Fu et al., 2024; Fei et al., 2024; Ye and Chen, 2024; Wang and Ma, 2024)
@@
Register (Wang et al., 2024g; Yang et al., 2024b)
Scan Flatten Bidirectional Scan (Zhu et al., 2024; Jiang et al., 2024b; Li and Chen, 2024; Li et al., 2024g)
Sweeping Scan (Liu et al., 2024c; Wang et al., 2024h; Yue and Li, 2024)
@@
Memory Initialization (Ezoe and Sato, 2024), Compression (Long et al., 2024; Nawrot et al., 2024), Connection (He et al., 2024c; Ren et al., 2024b)
@@ 4.1.1. Integration
  • Transformer-based models have exhibited remarkable performance in numerous tasks, but their quadratic computational complexity still hampers them during inference process (Gu et al., 2021a). In the pursuit of efficient generation, some researchers have proposed incorporating Mamba blocks with Transformer-based models. For example, Jamba (Lieber et al., 2024) combines blocks of Transformer and Mamba layers to tackle long-content Natural Language Processing tasks, capitalizing on the advantages of both model families. The Attention-Mamba hybrid model demonstrated superior performance compared to the standalone Transformer and Mamba models, achieving better throughput than the vanilla Transformer model. Mambaformer (Xu et al., 2024a) utilizes the hybrid framework to forecast multiple time series, including exchange rates, hourly electricity consumption, and power load, which internally combines Mamba blocks and Transformer layers for long- and short-range dependencies, respectively. Due to the integration of Mamba and Transformer, Mambaformer outperforms Transformer-based predictors in long-short range time series forecasting.
  • GNN has demonstrated promising potential in capturing neighboring relationships through message-passing mechanisms, where information is propagated over a connection graph through stacked layers. Nonetheless, these models face a significant limitation known as over-smoothing (Chen et al., 2020), particularly when attempting to capture high-order adjacency signals. To tackle such a challenge, Mamba has been employed for graph representation learning (Liu et al., 2024a; Li et al., 2024f; Yang et al., 2024d; Wang et al., 2024e). For example, Graph Mamba (Behrouz and Hashemi, 2024) reformulates graph-structured data into sequential tokens in a particular order and leverages a selective SSM layer within the Mamba block to construct a novel Graph Mamba Network (GMN) architecture, which achieves superior graph representation learning capabilities, particularly in the datasets that require high-order dependencies between nodes.
@@ 4.1.2. Substitution
  • U-Net. Many efforts (Shi et al., 2024b; Wang et al., 2024g, a; Liao et al., 2024) have been made to synergize U-Net with Mamba’s capability in capturing intricate and broad semantics so as to advance model performance in computer vision tasks. For example, Mamba-UNet (Wang et al., 2024g) utilizes Visual Mamba blocks exclusively to construct a U-Net-like model (i.e., an encoder-decoder model infused with skip connections) for medical image segmentation. Their evaluation demonstrates that Mamba-UNet surpasses several U-Net variations, which can be attributed to the efficacy and efficiency of Mamba blocks in handling long-range patch sequences.
  • Diffusion Model. Some endeavors (Fu et al., 2024; Fei et al., 2024; Oshima et al., 2024) have been undertaken to build a novel type of diffusion model, Diffusion State Space Model (DiS), which replace the typical backbone (e.g., CNNs, Attentions, U-Nets) with a state space backbone. Given the remarkable efficiency and efficacy of Mamba blocks in accommodating long-range dependencies, DiS is distinguished by generating longer sequences using diffusion models (Fei et al., 2024). For example, Oshima et al. (2024) propose a Mamba-based diffusion model that substantially decreases memory consumption for long video sequences, while still maintaining competitive performance metrics when compared to Transformer-based models. Moreover, MD-Dose (Fu et al., 2024) and P-Mamba (Ye and Chen, 2024) construct noise predictors using Mamba blocks in the backward process of diffusion models, ultimately generating specific targets for medical image processing.
@@ 4.2.1. Flattening Scan
  • Sweeping Scan. As illustrated in Figure 6, the Sweeping Scan technique processes the model inputs in a specific direction, analogous to a cleaner meticulously sweeping a floor (Yue and Li, 2024; Wang et al., 2024h). For instance, Cross-Scan (Liu et al., 2024c) entails dividing the input image into patches and subsequently flattening it along four distinct paths, which is regarded as a fusion of two bidirectional scans. By adopting these complementary traversal paths, Cross Scan enables each patch in the image to efficiently integrate information from its neighbors in different directions, thereby facilitating the establishment of informative, receptive fields. Omni-Scan (Shi et al., 2024b; Zhao et al., 2024a) incorporates the modeling of image information flows from multiple directions, e.g., 2 (forward and backward) $\times$ 4 (left-right, top-bottom, top right-bottom left, top left-bottom right). Such a strategy augments the global modeling capability of contextual information in various directions, enabling the extraction of comprehensive global spatial features.
@@ 4.3. Memory Management
Like RNNs, the memory of hidden states within state space models effectively stores information from previous steps, thereby playing a crucial role in SSM’s overall functionality. While Mamba has introduced the HiPPO-based method for memory initialization (Gu and Dao, 2023), challenges still exist in the memory management of the SSM cell, including transferring hidden information between layers and achieving lossless memory compression. To this end, a handful of pioneering studies have proposed different solutions, including memory initialization, compression, and connection. For example, Ezoe and Sato (2024) have attempted to refine the initialization process of selective SSMs by using a balanced truncation method during model retraining. Moreover, DGMamba (Long et al., 2024) introduces a Hidden State Suppressing method to bolster the domain generalization capabilities of the hidden states within State Space Models. This method works to alleviate the negative effects stemming from these hidden states, thereby narrowing the gap between hidden states across different domains. On a similar note, DenseMamba (He et al., 2024c) has put forth a dense connection method to enhance the propagation of hidden information between layers in SSMs. This strategy aims to mitigate memory degradation and preserve detailed information for output generation by selectively integrating hidden states from shallower layers into deeper ones.
@@ 5. Adapting Mamba to Diverse Data
Motion Continuous Human Motion Understanding (Wang et al., 2024d; Zhang et al., 2024c; Zeng et al., 2024)
Non-Sequential Data Image High-Resolution Medical Vision (Yue and Li, 2024; Chen et al., 2024a, d; Lin et al., 2024b; Wang et al., 2024g, g; Ruan and Xiang, 2024; Zhu et al., 2024; Touvron et al., 2021)
Graph Large Graph Learning (Fan et al., 2019b; Huang et al., 2020; Liu et al., 2024a; Ye and Ji, 2021; Wang et al., 2024e; Behrouz and Hashemi, 2024; Huang et al., 2024a)
@@
Multimodal Data Vision-Languge Visual and Linguistic Awareness (Yang et al., 2024c; Liu et al., 2024b; Wu et al., 2023; Wang et al., 2024d; Qiao et al., 2024)
@@ 5.1.1. Natural Language
As one of the most representative architectures, Mamba performs content-based reasoning while ensuring efficiency, which is considered a promising alternative for the backbone of large language models to address Transformers’ computational inefficiency on long sequences. Building upon this insight, many studies have explored the potential of Mamba for various downstream tasks in natural language processing (NLP) (Behrouz et al., 2024; Bhirangi et al., 2024; Nawrot et al., 2024; He et al., 2024c). For example, MambaByte (Wang et al., 2024b) is proposed to utilize Mamba on byte sequences, aiming to leverage the advantages of Mamba in capturing long-range dependencies for token-free language models. Their evaluations show that MambaByte avoids the inductive bias of subword tokenization and outperforms state-of-the-art subword Transformers on long-term language modeling tasks. Besides, Jamba (Lieber et al., 2024) and BlackMamba (Anthony et al., 2024) incorporate the concept of Mix-of-Experts (MoE) to enhance Mamba’s performance on language processing by integrating the linear-complexity generation from SSMs with the rapid and economical inference capabilities offered by MoE.
@@ 5.1.5. Motion
Human motion understanding and generation stand as a significant pursuit in a broad range of practical applications, including computer animation, game development, and robot manipulation. However, semantic actions that occur infrequently within lengthy motion sequences make long-range motion modeling difficult. To address this issue, several studies have proposed the use of Mamba to capture spatiotemporal patterns in motion sequences (Wang et al., 2024d). For instance, Motion Mamba (Zhang et al., 2024c) proposes a hybrid Mamba model, which leverages a hierarchical SSM layer to capture temporal patterns and introduces a bidirectional SSM layer to learn spatial knowledge, preserving motion consistency between frames. Based on the comprehensive experiments, the Mamba-based model outperforms representative diffusion-based methods in human motion generation tasks, achieving a 50% FID improvement and four times faster performance. Additionally, MambaMOS (Zeng et al., 2024) designs a motion-aware state space model that focuses explicitly on capturing variations in motion between consecutive time steps, which further emphasizes the exceptional capabilities of Mamba in achieving high-quality, lengthy sequence motion modeling.
@@ 5.2.1. Image
As one of the most prevalent modalities, image data forms the foundation of various computer vision applications, e.g., face recognition, medical vision (Yue and Li, 2024), and remote sensing (Chen et al., 2024a, d). Drawing inspiration from the success of Mamba in sequence modeling, there exists an intriguing opportunity to transfer this accomplishment from text processing to image analysis. It involves treating an image as a series of patches, potentially paving the way for new avenues of exploration within the realm of computer vision. Thus, plenty of Mamba-based vision models have recently been developed to alleviate heavy computational resources and memory pressures while exhibiting competitive modeling capabilities (Lin et al., 2024b; Wang et al., 2024g, g; Ruan and Xiang, 2024). For example, Vision Mamba (Zhu et al., 2024) incorporates bidirectional SSM to facilitate global visual semantic modeling and incorporates positional embeddings for location-aware visual comprehension. Not requiring attention mechanisms, Vision Mamba matches the modeling capacity of Vision Transformers while substantially decreasing computation time to subquadratic levels and upholding linear memory complexity. Specifically, it outperforms the state-of-the-art baseline DeiT (Touvron et al., 2021) in terms of speed, being 2.8× faster, and also presents a remarkable reduction of 86.8% in GPU memory usage during batch inference for feature extraction on high-resolution images (1248×1248). Moreover, VMamba (Liu et al., 2024c) introduces 2D Selective Scan (SS2D) that serves as a bridge between 1D array scanning and 2D plane traversal, enabling Mamba to process visual data effectively.

### v4

@@ 1. Introduction
Recently, a promising architecture, structured state space sequence models (SSMs) (Gu et al., 2021a), have emerged to efficiently capture complex dependencies in sequential data, becoming a formidable competitor to Transformer. These models, inspired by classical state space models (Kalman, 1960), can be considered a fusion of recurrent neural networks and convolutional neural networks. They can be computed efficiently using either recurrence or convolution operations, achieving linear or near-linear scaling with sequence length, thus significantly reducing the computational costs. More specifically, as one of the most successful SSM variants, Mamba achieves comparable modeling capabilities to Transformers while maintaining linear scalability with sequence length (Gu and Dao, 2023), propelling it into the realm of focal topics. Mamba first introduces a simple yet effective selection mechanism that enables the model to filter out irrelevant information while retaining necessary and relevant data indefinitely by parameterizing the SSM parameters based on the input. Then, Mamba proposes a hardware-aware algorithm to compute the model recurrently with a scan instead of convolution, achieving up to 3$\times$×faster computation on A100 GPUs. As shown in Figure 1, the powerful modeling capabilities for complex and lengthy sequential data, along with near-linear scalability, position Mamba as an emerging foundation model, poised to revolutionize various domains of research and applications, such as computer vision (Xu et al., 2024c; Zhu et al., 2024), natural language processing (Lieber et al., 2024; Zhao et al., 2024c), healthcare (Ruan and Xiang, 2024; Xing et al., 2024; Wang et al., 2024i), etc. For example, Zhu et al. (2024) propose Vim, which is 2.8$\times$×faster than DeiT (Touvron et al., 2021) and saves 86.8% GPU memory when extracting features for high-resolution images. Dao and Gu (2024) show the connections between SSMs and variants of attention and propose a new architecture that refines selective SSM, achieving 2-8$\times$ faster on language modeling.
@@
Concurrent with our survey, several related surveys have been released, purely focusing on state space models (Patro and Agneeswaran, 2024; Wang et al., 2024h) and Vision Mamba (Zhang et al., 2024h; Liu et al., 2024e; Xu et al., 2024c). Diverging from these surveys, this paper is centered on the associated research concerning Mamba. It systematically analyzes existing literature from a novel standpoint to explore the evolution of Mamba architecture and the data adaptation methods utilized in Mamba-based models.
@@ 3.2. Mamba-2: State Space Duality
(24) $\mathbf{y}=\textbf{SSD}(\mathbf{A},\mathbf{B},\mathbf{C})(\mathbf{x})=\mathbf{M}\mathbf{x},$
@@ 4. Advancements in Mamba Models
Block Integration Transformer (Lieber et al., 2024; Xu et al., 2024a; Pilault et al., 2024; Hatamizadeh and Kautz, 2024; Pitorro et al., 2024; Gao et al., 2024b; Park et al., 2024)
@@
Graph Neural Network (GNN) (Liu et al., 2024a; Li et al., 2024f; Behrouz and Hashemi, 2024; Wang et al., 2024g; Yang et al., 2024d)
@@
Substitution U-Net (Sepehri et al., 2024; Shi et al., 2024b; Wang et al., 2024i, a; Liu et al., 2024d; Ruan and Xiang, 2024; Liao et al., 2024; Ma and Wang, 2024; Sanjid et al., 2024; Deng and Gu, 2024; Ji et al., 2024; Hosseini et al., 2024)
Diffusion Models (Oshima et al., 2024; Fu et al., 2024a; Fei et al., 2024; Ye and Chen, 2024; Wang and Ma, 2024)
@@
Register (Wang et al., 2024i; Yang et al., 2024b)
Scan Flatten Bidirectional Scan (Zhu et al., 2024; Jiang et al., 2024b; Li and Chen, 2024; Li et al., 2024g; Qu et al., 2024c)
Sweeping Scan (Liu et al., 2024c; Wang et al., 2024j; Yue and Li, 2024)
@@
Memory Initialization (Ezoe and Sato, 2024), Compression (Long et al., 2024; Nawrot et al., 2024), Connection (He et al., 2024c; Ren et al., 2024b), State-Tracking (Merrill et al., 2024)
@@ 4.1.1. Integration
  • Transformer-based models have exhibited remarkable performance in numerous tasks, but their quadratic computational complexity still hampers them during inference process (Gu et al., 2021a). In the pursuit of efficient generation, some researchers have proposed incorporating Mamba blocks with Transformer-based models (Pilault et al., 2024; Hatamizadeh and Kautz, 2024; Pitorro et al., 2024; Gao et al., 2024b; Park et al., 2024). For example, Jamba (Lieber et al., 2024) combines blocks of Transformer and Mamba layers to tackle long-content Natural Language Processing tasks, capitalizing on the advantages of both model families. The Attention-Mamba hybrid model demonstrated superior performance compared to the standalone Transformer and Mamba models, achieving better throughput than the vanilla Transformer model. Mambaformer (Xu et al., 2024a) utilizes the hybrid framework to forecast multiple time series, which internally combines Mamba blocks and Transformer layers for long- and short-range dependencies, respectively. Due to the integration of Mamba and Transformer, Mambaformer outperforms Transformer-based predictors in long-short range time series forecasting.
  • GNN has demonstrated promising potential in capturing neighboring relationships through message-passing mechanisms, where information is propagated over a connection graph through stacked layers. Nonetheless, these models face a significant limitation known as over-smoothing (Chen et al., 2020), particularly when attempting to capture high-order adjacency signals. To tackle such a challenge, Mamba has been employed for graph representation learning (Liu et al., 2024a; Li et al., 2024f; Yang et al., 2024d; Wang et al., 2024g). For example, Graph Mamba (Behrouz and Hashemi, 2024) reformulates graph-structured data into sequential tokens in a particular order and leverages a selective SSM layer within the Mamba block to construct a novel Graph Mamba Network (GMN) architecture, which achieves superior graph representation learning capabilities, particularly in the datasets that require high-order dependencies between nodes.
@@ 4.1.2. Substitution
  • U-Net. Many efforts (Shi et al., 2024b; Wang et al., 2024i, a; Liao et al., 2024) have been made to synergize U-Net with Mamba’s capability in capturing intricate and broad semantics so as to advance model performance in computer vision tasks. For example, Mamba-UNet (Wang et al., 2024i) utilizes Visual Mamba blocks exclusively to construct a U-Net-like model (i.e., an encoder-decoder model infused with skip connections) for medical image segmentation. Their evaluation demonstrates that Mamba-UNet surpasses several U-Net variations, which can be attributed to the efficacy and efficiency of Mamba blocks in handling long-range patch sequences.
  • Diffusion Model. Some endeavors (Fu et al., 2024a; Fei et al., 2024; Oshima et al., 2024) have been undertaken to build a novel type of diffusion model, Diffusion State Space Model (DiS), which replace the typical backbone (e.g., CNNs, Attentions, U-Nets) with a state space backbone. Given the remarkable efficiency and efficacy of Mamba blocks in accommodating long-range dependencies, DiS is distinguished by generating longer sequences using diffusion models (Fei et al., 2024). For example, Oshima et al. (2024) propose a Mamba-based diffusion model that substantially decreases memory consumption for long video sequences, while still maintaining competitive performance metrics when compared to Transformer-based models. Moreover, MD-Dose (Fu et al., 2024a) and P-Mamba (Ye and Chen, 2024) construct noise predictors using Mamba blocks in the backward process of diffusion models, ultimately generating specific targets for medical image processing.
@@ 4.2.1. Flattening Scan
  • Sweeping Scan. As illustrated in Figure 6, the Sweeping Scan technique processes the model inputs in a specific direction, analogous to a cleaner meticulously sweeping a floor (Yue and Li, 2024; Wang et al., 2024j). For instance, Cross-Scan (Liu et al., 2024c) entails dividing the input image into patches and subsequently flattening it along four distinct paths, which is regarded as a fusion of two bidirectional scans. By adopting these complementary traversal paths, Cross Scan enables each patch in the image to efficiently integrate information from its neighbors in different directions, thereby facilitating the establishment of informative, receptive fields. Omni-Scan (Shi et al., 2024b; Zhao et al., 2024a) incorporates the modeling of image information flows from multiple directions, e.g., 2 (forward and backward) $\times$ 4 (left-right, top-bottom, top right-bottom left, top left-bottom right). Such a strategy augments the global modeling capability of contextual information in various directions, enabling the extraction of comprehensive global spatial features.
@@ 4.3. Memory Management
Like RNNs, the memory of hidden states within state space models effectively stores information from previous steps, thereby playing a crucial role in SSM’s overall functionality. While Mamba has introduced the HiPPO-based method for memory initialization (Gu and Dao, 2023), challenges still exist in the memory management of the SSM cell, including transferring hidden information between layers and achieving lossless memory compression. To this end, a handful of pioneering studies have proposed different solutions (Wang et al., 2024e). For example, Ezoe and Sato (2024) have attempted to refine the initialization process of selective SSMs by using a balanced truncation method during model retraining. Moreover, DGMamba (Long et al., 2024) introduces a Hidden State Suppressing method to bolster the domain generalization capabilities of the hidden states within State Space Models. This method works to alleviate the negative effects stemming from these hidden states, thereby narrowing the gap between hidden states across different domains. On a similar note, DenseMamba (He et al., 2024c) has put forth a dense connection method to enhance the propagation of hidden information between layers in SSMs. This strategy aims to mitigate memory degradation and preserve detailed information for output generation by selectively integrating hidden states from shallower layers into deeper ones. Moreover, Merrill et al. (2024) delve into the concept of the illusion of ”State” in SSMs through extensive experiments, revealing their constraints in addressing practical state-tracking challenges like monitoring chess moves. Expanding on this, the researchers enhance SSMs with a subtle modification: by introducing input-dependent transition matrices, enabling effective state tracking and permutation composition.
@@ 5. Adapting Mamba to Diverse Data
Motion Continuous Human Motion Understanding (Wang et al., 2024f; Zhang et al., 2024c; Zeng et al., 2024)
Non-Sequential Data Image High-Resolution Medical Vision (Yue and Li, 2024; Chen et al., 2024a, d; Lin et al., 2024b; Wang et al., 2024i, i; Ruan and Xiang, 2024; Zhu et al., 2024; Touvron et al., 2021)
Graph Large Graph Learning (Fan et al., 2019b; Huang et al., 2020; Liu et al., 2024a; Ye and Ji, 2021; Wang et al., 2024g; Behrouz and Hashemi, 2024; Huang et al., 2024a)
@@
Multimodal Data Vision-Languge Visual and Linguistic Awareness (Yang et al., 2024c; Liu et al., 2024b; Wu et al., 2023; Wang et al., 2024f; Qiao et al., 2024)
@@ 5.1.1. Natural Language
As one of the most representative architectures, Mamba performs content-based reasoning while ensuring efficiency, which is considered a promising alternative for the backbone of large language models to address Transformers’ computational inefficiency on long sequences. Building upon this insight, many studies have explored the potential of Mamba for various downstream tasks in natural language processing (NLP) (Behrouz et al., 2024; Bhirangi et al., 2024; Nawrot et al., 2024; He et al., 2024c). For example, MambaByte (Wang et al., 2024c) is proposed to utilize Mamba on byte sequences, aiming to leverage the advantages of Mamba in capturing long-range dependencies for token-free language models. Their evaluations show that MambaByte avoids the inductive bias of subword tokenization and outperforms state-of-the-art subword Transformers on long-term language modeling tasks. Besides, Jamba (Lieber et al., 2024) and BlackMamba (Anthony et al., 2024) incorporate the concept of Mix-of-Experts (MoE) to enhance Mamba’s performance on language processing by integrating the linear-complexity generation from SSMs with the rapid and economical inference capabilities offered by MoE.
@@ 5.1.5. Motion
Human motion understanding and generation stand as a significant pursuit in a broad range of practical applications, including computer animation, game development, and robot manipulation. However, semantic actions that occur infrequently within lengthy motion sequences make long-range motion modeling difficult. To address this issue, several studies have proposed the use of Mamba to capture spatiotemporal patterns in motion sequences (Wang et al., 2024f). For instance, Motion Mamba (Zhang et al., 2024c) proposes a hybrid Mamba model, which leverages a hierarchical SSM layer to capture temporal patterns and introduces a bidirectional SSM layer to learn spatial knowledge, preserving motion consistency between frames. Based on the comprehensive experiments, the Mamba-based model outperforms representative diffusion-based methods in human motion generation tasks, achieving a 50% FID improvement and four times faster performance. Additionally, MambaMOS (Zeng et al., 2024) designs a motion-aware state space model that focuses explicitly on capturing variations in motion between consecutive time steps, which further emphasizes the exceptional capabilities of Mamba in achieving high-quality, lengthy sequence motion modeling.
@@ 5.2.1. Image
As one of the most prevalent modalities, image data forms the foundation of various computer vision applications, e.g., face recognition, medical vision (Yue and Li, 2024), and remote sensing (Chen et al., 2024a, d). Drawing inspiration from the success of Mamba in sequence modeling, there exists an intriguing opportunity to transfer this accomplishment from text processing to image analysis. It involves treating an image as a series of patches, potentially paving the way for new avenues of exploration within the realm of computer vision. Thus, plenty of Mamba-based vision models have recently been developed to alleviate heavy computational resources and memory pressures while exhibiting competitive modeling capabilities (Lin et al., 2024b; Wang et al., 2024i, i; Ruan and Xiang, 2024; Fu et al., 2024b; Wang et al., 2024b). For example, Vision Mamba (Zhu et al., 2024) incorporates bidirectional SSM to facilitate global visual semantic modeling and incorporates positional embeddings for location-aware visual comprehension. Not requiring attention mechanisms, Vision Mamba matches the modeling capacity of Vision Transformers while substantially decreasing computation time to subquadratic levels and upholding linear memory complexity. Specifically, it outperforms the state-of-the-art baseline DeiT (Touvron et al., 2021) in terms of speed, being 2.8× faster, and also presents a remarkable reduction of 86.8% in GPU memory usage during batch inference for feature extraction on high-resolution images (1248×1248). Moreover, VMamba (Liu et al., 2024c) introduces 2D Selective Scan (SS2D) that serves as a bridge between 1D array scanning and 2D plane traversal, enabling Mamba to process visual data effectively. Additionally, LKM-UNet (Wang et al., 2024b) presents a Mamba-based model designed for medical image segmentation. This model leverages large Mamba kernels within a UNet architecture, showcasing superior performance in both local and global spatial modeling for visual inputs in medical imaging tasks.